OpenStudy (anonymous):

help please? Medal & Fan I am not looking for the answer to anything, I just need someone to help me learn how to find the volume of solid figures. (p.s. idk what 'volume' is eiter)

OpenStudy (anonymous):

nvm XD

OpenStudy (anonymous):

ok volume of wat

OpenStudy (anonymous):

of a cone

OpenStudy (anonymous):

\[\pi r^2 h/3\]

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

\[\pi \] =3.14

OpenStudy (anonymous):

okay, so I have this math problem. A cone is inscribed in a regular square pyramid. If the pyramid has a base edge of 6" and a slant height of 9", find the volume of the cone. how would I use this information to fill out that formula. sorry, i am really bad at math. i need like a full talk through, if u can give me one

OpenStudy (anonymous):

\[r^2\] = half the length of the bottom circle squared

OpenStudy (anonymous):

wait for like 5 minutes

OpenStudy (anonymous):

k :)

OpenStudy (anonymous):

O_O my bro has the same question for his homework ima copy and paste XD

OpenStudy (anonymous):

area of the cone will equal 1/3 (area of the base) (height) for the area of the base, we know that the cone has a diameter of 6" (since it is the largest possible circle inside the square base). so use the area of a circle formula to get: (pi)r^2 where r=3 (1/2 of 6) 9(pi) the height is a little tickier, since you know the slant height, you need to use the properties of a right triangle to get the height. The hypotenuse is 9" and the base is 3" so we use a^2 + b^2 = c^2 3^2 + b^2= 9^2 9+b^2=81 b^2=72 b=8.49 now that you have the height and the area of the base, put it all together to get: 1/3 (9(pi))(8.49) your answer is 25.47 in^3

OpenStudy (anonymous):

okay, how you did it makes sense to me, but the choices are 12sqrt 2pie 18sqrt 2pie 27pie

OpenStudy (anonymous):

well that would be easier

OpenStudy (anonymous):

u still there?

OpenStudy (anonymous):

ya :c im confused

OpenStudy (anonymous):

b

OpenStudy (anonymous):

thnx but can you explain how u got it? :P

OpenStudy (anonymous):

nvm i gov up kill me

OpenStudy (anonymous):

@xapproachesinfinity help meh bro

OpenStudy (anonymous):

lol k thanks tho XD have a good night :)

OpenStudy (anonymous):

@xapproachesinfinity u there???

OpenStudy (anonymous):

@perl could u help?

OpenStudy (amorfide):

@rational